# Skewness-dependent Moments of Pion GPD from Nonlocal Quark-Bilinear Correlators

### v1

@@ I Introduction
Experimentally, GPDs are accessed primarily through exclusive processes such as deeply virtual Compton scattering (DVCS) [8] and deeply virtual meson production (DVMP) [9, 10]. Over the past two decades, extensive DVCS and DVMP measurements on the nucleon have provided valuable constraints on nucleon GPDs [11, 12, 13, 14, 15, 16], although their extraction remains challenging due to convolution integrals in the observables [17, 18, 19, 20]. In contrast, experimental information on pion GPDs is still very limited. Current knowledge is restricted to related quantities such as the pion electromagnetic form factor [21, 22] and PDFs from Drell–Yan processes [23, 24, 25, 26]. Direct constraints on pion GPDs are absent, but upcoming programs at Jefferson Lab [27], AMBER at CERN [28], the Electron–Ion Collider (EIC) [29], and the proposed Electron–Ion Collider in China (EicC) [30] are expected to provide valuable information. In this context, lattice QCD plays a crucial role by supplying first-principles input on pion GPDs and guiding future experimental efforts.
@@
With the advent of large-momentum effective theory (LaMET) [69, 70, 71], it has become possible to calculate the $x$-dependence of parton distributions, including the GPDs [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91], from lattice QCD. The central idea is to evaluate matrix elements of equal-time nonlocal quark-bilinear operators with external hadrons boosted to large momenta and relate them to light-cone quantities through large-momentum expansion and perturbative matching. The same correlators also provide complementary access to Mellin moments through the pseudo-distribution or short-distance factorization (SDF) approach [92, 93], which allows for a model-independent determination of moments through OPE [94]. This strategy has been applied to pion and proton PDFs [95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109], light-cone distribution amplitudes (DAs) of mesons [110, 111, 112], and, more recently, extended to nucleon and pion GPDs [113, 114, 115], where higher moments can be extracted without the power-divergent operator mixing that limits the local-operator approach.
@@
The rest of the paper is organized as follows: In Section II, we introduce the definition of pion quasi-GPD in Euclidean space and outline the methodology for extracting the relevant matrix elements. We also discuss the renormalization, perturbative matching, combined fits, as well as logarithmic renormalization-group resummation (RGR). Section III presents our lattice setup and describes the extraction of ground-state matrix elements from two- and three-point correlation functions. In Section IV, we present our simulation results, including the pion GPD moments and the final combined fit results across different momentum transfers and skewness values. We compare analyses at fixed order with those including RGR at NLL accuracy. Finally, Section V summarizes our findings.
@@ II.1 Quasi-GPD matrix elements on the lattice
$H^{\pi}(x,\xi,t)$ $=\int\frac{dz^{-}}{4\pi}e^{ixP^{+}z^{-}}\left\langle\pi(P_{f})\left|\bar{\psi}\left(-\frac{z^{-}}{2}\right)\gamma^{+}\mathcal{L}\left(-\frac{z^{-}}{2},\frac{z^{-}}{2}\right)\psi\left(\frac{z^{-}}{2}\right)\right|\pi(P_{i})\right\rangle,$ (1)
@@
$\mathcal{O}^{\mu}=\frac{1}{2}\left[\bar{u}\left(-\frac{z}{2}\right)\gamma^{\mu}\mathcal{L}\left(-\frac{z}{2},\frac{z}{2}\right)u\left(\frac{z}{2}\right)-\bar{d}\left(-\frac{z}{2}\right)\gamma^{\mu}\mathcal{L}\left(-\frac{z}{2},\frac{z}{2}\right)d\left(\frac{z}{2}\right)\right],$ (3)
where $z$ is the component of $z^{\mu}=(0,0,0,z)$ in the spacial-direction. The isovector construction eliminates all quark-disconnected contributions by treating the $u$ and $d$ quarks as degenerate under isospin symmetry. In this limit, the isovector pion GPD coincides with the valence-quark GPD [100]. Furthermore, G-parity dictates that the $u$- and $d$-quark bilinears in Eq. (3) differ only by the sign of their real part while sharing the same imaginary part, so their difference yields a purely real contribution [122]. In practical lattice calculations, this means that one needs only to evaluate the $u$-quark operator and take its real part.
@@
$\widetilde{\mathcal{H}}(z,P_{z},\xi,t)=\mathcal{A}_{1}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2})+\frac{\Delta^{0}}{P^{0}}\,\mathcal{A}_{3}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2}).$ (7)
@@
$\widetilde{\mathcal{H}}_{\text{LI}}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2})=\mathcal{A}_{1}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2})+\frac{z\cdot\Delta}{z\cdot P}\,\mathcal{A}_{3}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2}),$ (8)
@@ II.2 Renormalization and short distance factorization
$\widetilde{\mathcal{H}}^{R}(z,P_{z},\xi,t)=\sum_{n=1}^{\infty}\frac{(-izP_{z})^{n-1}}{{(n-1)}!}\sum_{k=1}^{n}H_{k}(\xi,t,\mu)\,\xi^{\,n-k}\,c_{n,n-k}(z^{2}\mu^{2})+\mathcal{O}(\Lambda_{\mathrm{QCD}}^{2}z^{2}),$ (10)
@@
Here $A_{n,k}(t)$ and $C_{n}(t)$ denote the tower of GFFs. Eq. (12) makes polynomiality explicit; it provides a fundamental constraint on the behavior of GPDs, informs consistent parameterizations in global analyses, and offers a stringent internal consistency check for lattice determinations of Mellin moments.
@@ II.3 Implementation of the RG resummation
Resummation order $\alpha^{n}L^{k}$ log resummed Anomalous dimension $\beta$-funtion $c_{n,n-k}$
@@
Table 2: Summary of the logarithmic accuracies used in the RG-improved $c_{n,n-k}$. The corresponding loop orders of the anomalous dimension, $\beta$-function, and the matching coefficient are listed for each level.
@@
To achieve NNLL accuracy, the anomalous dimension and the QCD $\beta$–function to be known through three loops,
@@ III.1 Lattice Setup
We employ the same lattice setup as in our previous work [87], utilizing 2+1 flavor Highly Improved Staggered Quark (HISQ) [135] gauge ensembles generated by the HotQCD collaboration [136] with a lattice spacing of $a=0.04$ fm on a $64^{3}\times 64$ grid. The sea quark masses correspond to a pion mass of 160 MeV. In the valence sector, we use the Wilson-Clover action with one level of HYP smearing [137], and the clover coefficient is set to the tree-level tadpole-improved value, 1.02868 [100]. The valence quark mass is tuned to $am_{q}=-0.033$, yielding a pion mass of 300 MeV.
@@ III.2 Bare quasi-GPD Matrix Elements
$C_{{\rm 3pt}}(\mathbf{P}^{f},\mathbf{q};t_{s},\tau;z)=\sum_{\mathbf{x,y}}e^{-i\mathbf{P}^{f}\cdot\mathbf{x}}e^{i\mathbf{q}\cdot\mathbf{y}}\left\langle\pi_{\rm s}(\mathbf{x},t_{s})\mathcal{O}^{\mu}(\mathbf{y},\tau;z)\pi_{\rm s}^{\dagger}(\mathbf{0},0)\right\rangle,$ (22)
where $\mathbf{P}=2\pi\mathbf{n}/(aN_{s})$ denotes the spatial momentum, with $\mathbf{n}$ being the dimensionless lattice unit. Here, $\mathbf{q}=\mathbf{P}^{f}-\mathbf{P}^{i}$ is the spatial momentum difference between the final and initial states. The variables $t_{s}$ represent the temporal separations between the final and initial states, respectively. Additionally, the operators $\pi^{\dagger}_{s}$, $\pi_{s}$, and $\mathcal{O}^{\mu}$ stand for the pion creation, annihilation operators, and the insertion operator with component $\gamma^{\mu}$, respectively. The subscripts $\rm s$}” indicates “smeared” fields. The current operator $\mathcal{O}^{\mu}$, defined in Eq. (3), is inserted at spacetime position ($\mathbf{y},\tau$) and contains a straight Wilson-line of length $z$.
@@
$R^{fi}(\mathbf{P}^{f},\mathbf{P}^{i};t_{s},\tau;z)\!\equiv\!\frac{2\sqrt{E_{0}^{f}E_{0}^{i}}}{E_{0}^{f}+E_{0}^{i}}\frac{C_{{\rm 3pt}}(\mathbf{P}^{f},\mathbf{P^{i}};t_{s},\tau;z)}{C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s})}\left[\frac{C_{{\rm 2pt}}(\mathbf{P}^{i},t_{s}\!-\!\tau)C_{{\rm 2pt}}(\mathbf{P}^{f},\tau)C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s})}{C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s}\!-\!\tau)C_{{\rm 2pt}}(\mathbf{P}^{i},\tau)C_{{\rm 2pt}}(\mathbf{P}^{i},t_{s})}\right]^{1/2},$ (23)
@@
We then fit the lattice results for $R^{fi}$ using the corresponding spectral decomposition ansatz, as details in Ref. [145]. In Fig. 2, we present the fit results for the ratio $R^{fi}$ at two different non-zero skewness values. The left panel shows the case at $\xi=0.14$ with t=0.597-t=0.597 GeV2, while the right panel corresponds to a larger momentum transfer of t=2.748-t=2.748 GeV2 at $\xi=-0.33$. The lattice results are displayed as data points, with different colors indicating different source-sink time separations $t_{s}$. The colored bands represent the fit results for the corresponding datasets, which show good agreement with the lattice results. By taking the limits as $t_{s}\rightarrow\infty$ and $\tau\rightarrow\infty$, we extract the bare matrix element for the ground state, namely,
@@
Figure 3: Comparison of the quasi-GPDs in coordinate space defined via the LI construction ($\widetilde{\mathcal{H}}_{\rm LI}$) and the $\gamma_{0}$ operator ($\widetilde{\mathcal{H}}$). Left: $P_{z}=1.937$ GeV, t=0.864-t=0.864 GeV2, right: $P_{z}=1.937$ GeV, t=2.748-t=2.748 GeV2.
@@ III.3 Renormalized quasi-GPD Matrix Elements
To obtain physically meaningful results from lattice QCD calculations, the bare matrix elements extracted from correlation functions must be properly renormalized. In this work, we adopt the ratio scheme for renormalization as shown in Eq. (9). Fig. 4 shows the renormalized matrix elements $\widetilde{\mathcal{H}}^{R}$ as a function of $\lambda$ for four representative nonzero values of $\xi$. In each panel at fixed $\xi$, the magnitude of the matrix element decreases monotonically with increasing $|\lambda|$ for all momentum transfers, as expected from the suppression of long-distance contributions in the coordinate-space correlator. Additionally, at fixed $\lambda$ within each panel, we observe a systematic falloff in the matrix element magnitude as the momentum transfer t-t increases, consistent with the generic $t$-slope seen in hadronic form factors. These trends are consistent with short-distance dominance at small $z^{2}$ and provide stable inputs for the polynomiality-constrained fits employed in our moment extraction.
Figure 4: Renoramlized matrix elements $\widetilde{\mathcal{H}}^{R}$ as functions of $\lambda$ at four nonzero skewness values. In each panel, data points are distinguished by color, representing results at different momentum transfers t-t.
@@ IV.1 Fitting strategy
For the fitting range, we fix the lower bound at $z_{\min}=0.08$ fm and vary the upper bound $z_{\max}$ to determine an optimal interval. Since the short-distance factorization is applicable only at short distances, we restrict $z_{\max}$ to the range $0.20\text{--}0.44~\mathrm{fm}$ and compare results obtained with LL, NLL, and NNLL resummation. Fig. 5 illustrates an example of the extracted moments at momentum $P_{z}=1.453~\mathrm{GeV}$ and momentum transfer $t=0.446~\mathrm{GeV}^{2}$, with the renormalization scale fixed at $\mu=2~\mathrm{GeV}$. We vary $\kappa$ among $1$, $\sqrt{2}$, and $2$, which are still close to one, to estimate the associated uncertainty. This range was also chosen after a study of the short-distance behavior of the same zero-momentum matrix elements in Ref. [87]. For the lowest moment $A_{1,0}$, RGR effects are negligible because it is associated with the conserved vector current, whose anomalous dimension vanishes. This behavior is indeed observed in the first panel, where the fixed–order and resummed results are essentially indistinguishable. In contrast, $A_{3,0}$ has a nonzero anomalous dimension, so logarithmic $\ln(z^{2}\mu^{2})$ terms appear in the matching, leading to visible sensitivity to the resummation as the fit range $z_{\max}$ is varied. At larger $z_{\max}$, the physical scale $\mu_{0}={(2\kappa e^{-\gamma_{E}})}/{|z|}$ decreases, driving the running coupling $\alpha_{s}(\mu_{0})$ into the nonperturbative regime and thereby spoiling the reliability of perturbation theory. From the $A_{3,0}$ panels, we observe that the LL-resummed result (matching coefficients at tree level) lies systematically below the NLL– and NNLL–resummed ones, illustrating the impact of perturbative matching. Moreover, varying $\kappa=\{1,\sqrt{2},2\}$ shows that the plateau region becomes progressively broader as $\kappa$ increases. This is mainly because larger $\kappa$ allows for a wider range of $z$ to be included in the fit, while the weight is mainly carried by the short-$z$ matrix elements. To ensure stability and control systematic uncertainties, we therefore restrict the analysis to a short-distance region where fixed order (up to NNLO), NLL-, and NNLL-resummed results agree within the statistical uncertainties. Therefore, this is a definitive choice based on the criteria given, so it is not conservative; we adopt $z_{\max}=0.24~\mathrm{fm}$ in the following analysis. We further remark that when $\kappa\sim 1$ the nonperturbative effects are still sizable and do not provide meaningful constraints on $A_{5,0}$. In particular, no clear plateau is observed at $\kappa=1$, while the NNLL results show visibly better convergence than NLL. Accordingly, our choice of $z_{\max}$ is primarily informed by the plateau behavior observed in $A_{3,0}$.

### v2

# Skewness-dependent moments of the pion GPD from nonlocal quark-bilinear correlators

@@ I Introduction
Experimentally, GPDs are accessed primarily through exclusive processes such as deeply virtual Compton scattering (DVCS) [8] and deeply virtual meson production (DVMP) [9, 10]. Over the past two decades, extensive DVCS and DVMP measurements on the nucleon have provided valuable constraints on nucleon GPDs [11, 12, 13, 14, 15], although their extraction remains challenging due to convolution integrals in the observables [16, 17, 18, 19]. Beyond DVCS/DVMP, phenomenological fits have also used form-factor data to constrain GPD parameterizations for the nucleon [20] and the pion [21]. In contrast, experimental information on pion GPDs is still very limited. Current knowledge is restricted to related quantities such as the pion electromagnetic form factor [22, 23] and PDFs from Drell–Yan processes [24, 25, 26, 27]. Direct constraints on pion GPDs are absent, but upcoming programs at Jefferson Lab [28], AMBER at CERN [29], the Electron–Ion Collider (EIC) [30], and the proposed Electron–Ion Collider in China (EicC) [31] are expected to provide valuable information. In this context, lattice QCD plays a crucial role by supplying first-principles input on pion GPDs and guiding future experimental efforts.
@@
With the advent of large-momentum effective theory (LaMET) [70, 71, 72], it has become possible to calculate the $x$-dependence of parton distributions, including the GPDs [73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92], from lattice QCD. The central idea is to evaluate matrix elements of equal-time nonlocal quark-bilinear operators with external hadrons boosted to large momenta and relate them to light-cone quantities through large-momentum expansion and perturbative matching. The same correlators also provide complementary access to Mellin moments through the pseudo-distribution or short-distance factorization (SDF) approach [93, 94], which allows for a model-independent determination of moments through OPE [95]. This strategy has been applied to pion and proton PDFs [96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110], light-cone distribution amplitudes (DAs) of mesons [111, 112, 113, 114], and, more recently, extended to nucleon and pion GPDs [115, 116, 117], where higher moments can be extracted without the power-divergent operator mixing that limits the local-operator approach.
@@
The rest of the paper is organized as follows: In Sec. II, we introduce the definition of pion quasi-GPD in Euclidean space and outline the methodology for extracting the relevant matrix elements. We also discuss the renormalization, perturbative matching, combined fits, as well as logarithmic renormalization-group resummation (RGR). Section III presents our lattice setup and describes the extraction of ground-state matrix elements from two- and three-point correlation functions. In Sec. IV, we present our simulation results, including the pion GPD moments and the final combined fit results across different momentum transfers and skewness values. We compare analyses at fixed order with those including RGR at NLL accuracy. Finally, Sec. V summarizes our findings.
@@ II.1 Quasi-GPD matrix elements on the lattice
$H^{\pi}(x,\xi,t)=\int\frac{\mathrm{d}z^{-}}{4\pi}\,$ $e^{ixP^{+}z^{-}}\big\langle\pi(P_{f})\big|\bar{\psi}\!\left(-\tfrac{z^{-}}{2}\right)\gamma^{+}$
$\times\mathcal{L}\!\left(-\tfrac{z^{-}}{2},\tfrac{z^{-}}{2}\right)\psi\!\left(\tfrac{z^{-}}{2}\right)\big|\pi(P_{i})\big\rangle,$ (1)
@@
$\mathcal{O}^{\mu}=\frac{1}{2}$ $\bigg[\bar{u}\left(-\frac{z}{2}\right)\gamma^{\mu}\mathcal{L}\left(-\frac{z}{2},\frac{z}{2}\right)u\left(\frac{z}{2}\right)$
$-\bar{d}\left(-\frac{z}{2}\right)\gamma^{\mu}\mathcal{L}\left(-\frac{z}{2},\frac{z}{2}\right)d\left(\frac{z}{2}\right)\bigg],$ (3)
where $z$ is the component of $z^{\mu}=(0,0,0,z)$ in the spacial-direction. The isovector construction eliminates all quark-disconnected contributions by treating the $u$ and $d$ quarks as degenerate under isospin symmetry. In this limit, the isovector pion GPD coincides with the valence-quark GPD [101]. Furthermore, G-parity dictates that the $u$- and $d$-quark bilinears in Eq. (II.1) differ only by the sign of their real part while sharing the same imaginary part, so their difference yields a purely real contribution [124]. In practical lattice calculations, this means that one needs only to evaluate the $u$-quark operator and take its real part.
@@
$\widetilde{\mathcal{H}}$ $(z,P_{z},\xi,t)$
$=\mathcal{A}_{1}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2})+\frac{\Delta^{0}}{P^{0}}\,\mathcal{A}_{3}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2}).$ (7)
@@
$\widetilde{\mathcal{H}}_{\text{LI}}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2})$
$\quad\!=\!\mathcal{A}_{1}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2})\!+\!\frac{z\cdot\Delta}{z\cdot P}\mathcal{A}_{3}(z\cdot P,z\cdot\Delta,\Delta^{2},z^{2}),$ (8)
@@ II.2 Renormalization and short distance factorization
$\widetilde{\mathcal{H}}^{R}$ $(z,P_{z},\xi,t)$
$=\sum_{n=1}^{\infty}\frac{(-izP_{z})^{n-1}}{{(n-1)}!}\sum_{k=1}^{n}H_{k}(\xi,t,\mu)\,\xi^{\,n-k}\,c_{n,n-k}(z^{2}\mu^{2}),$ (10)
@@
Here $A_{n,k}(t)$ and $C_{n}(t)$ denote the tower of GFFs. Equation (12) makes polynomiality explicit; it provides a fundamental constraint on the behavior of GPDs, informs consistent parameterizations in global analyses, and offers a stringent internal consistency check for lattice determinations of Mellin moments.
@@ II.3 Implementation of the RG resummation
   Resummation order    $\alpha^{n}L^{k}$ log resummed    Anomalous dimension    $\beta$-function    $c_{n,n-k}$
@@
Table 2: Summary of the logarithmic accuracies used in the RG-improved $c_{n,n-k}$. The corresponding loop orders of the anomalous dimension, $\beta$-function, and the matching coefficients are listed for each level.
@@
To achieve NNLL accuracy, the anomalous dimension and the QCD $\beta$ function are required to be known through three loops:
@@ III.1 Lattice Setup
We employ the same lattice setup as in our previous work [88], utilizing 2+1 flavor Highly Improved Staggered Quark (HISQ) [137] gauge ensembles generated by the HotQCD collaboration [138] with a lattice spacing of $a=0.04$ fm on a $64^{3}\times 64$ grid. The sea quark masses correspond to a pion mass of 160 MeV. In the valence sector, we use the Wilson-Clover action with one level of HYP smearing [139], and the clover coefficient is set to the tree-level tadpole-improved value, 1.02868 [101]. The valence quark mass is tuned to $am_{q}=-0.033$, yielding a pion mass of 300 MeV. The valence pion mass is set to 300 MeV to maintain a manageable signal-to-noise ratio for boosted correlators, consistent with observations in our previous work [101, 140]; approaching the 160 MeV sea mass leads to prohibitively noisy matrix elements. Calculations at lighter valence masses will be pursued in future work.
@@ III.2 Bare quasi-GPD Matrix Elements
$C_{{\rm 3pt}}$ $(\mathbf{P}^{f},\mathbf{q};t_{s},\tau;z)$
$=\sum_{\mathbf{x,y}}e^{-i\mathbf{P}^{f}\cdot\mathbf{x}}e^{i\mathbf{q}\cdot\mathbf{y}}\left\langle\pi_{\rm s}(\mathbf{x},t_{s})\mathcal{O}^{\mu}(\mathbf{y},\tau;z)\pi_{\rm s}^{\dagger}(\mathbf{0},0)\right\rangle,$ (22)
where $\mathbf{P}=2\pi\mathbf{n}/(aN_{s})$ denotes the spatial momentum, with $\mathbf{n}$ being the dimensionless lattice unit. Here, $\mathbf{q}=\mathbf{P}^{f}-\mathbf{P}^{i}$ is the spatial momentum difference between the final and initial states. The variables $t_{s}$ represent the temporal separations between the final and initial states, respectively. Additionally, the operators $\pi^{\dagger}_{s}$, $\pi_{s}$, and $\mathcal{O}^{\mu}$ stand for the pion creation, annihilation operators, and the insertion operator with component $\gamma^{\mu}$, respectively. The subscripts $\rm s$}” indicates “smeared” fields. The current operator $\mathcal{O}^{\mu}$, defined in Eq. (II.1), is inserted at spacetime position ($\mathbf{y},\tau$) and contains a straight Wilson-line of length $z$.
@@
$R^{fi}$ $(\mathbf{P}^{f},\mathbf{P}^{i};t_{s},\tau;z)\!\equiv\!\frac{2\sqrt{E_{0}^{f}E_{0}^{i}}}{E_{0}^{f}+E_{0}^{i}}\frac{C_{{\rm 3pt}}(\mathbf{P}^{f},\mathbf{P^{i}};t_{s},\tau;z)}{C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s})}$
$\times\left[\frac{C_{{\rm 2pt}}(\mathbf{P}^{i},t_{s}\!-\!\tau)C_{{\rm 2pt}}(\mathbf{P}^{f},\tau)C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s})}{C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s}\!-\!\tau)C_{{\rm 2pt}}(\mathbf{P}^{i},\tau)C_{{\rm 2pt}}(\mathbf{P}^{i},t_{s})}\right]^{1/2},$ (23)
@@
In Fig. 2, we present the ratio $R^{fi}$ for two representative non-zero skewness values. The upper panel shows the case of $\xi=-0.14$ with t=0.597-t=0.597 GeV2, while the lower panel corresponds to a larger momentum transfer of t=2.748-t=2.748 GeV2 at $\xi=-0.33$. In each panel, the data points correspond to the lattice results for three source-sink time separations, $t_{s}/a\in\{9,12,15\}$. The visible curvature in the lattice results indicates the presence of excited-state contamination. To control these effects, we employed a two-state fit analysis utilizing multiple time separations to isolate the ground-state matrix elements (see Ref. [148] for methodological details). The colored bands represent the resulting fit curves, which show good agreement with the lattice results. Finally, by extrapolating these fit results to the infinite time separation as $t_{s}\rightarrow\infty$ and $\tau\rightarrow\infty$, one can extract the bare matrix elements for the ground state, namely,
@@
Figure 3: Comparison of the quasi-GPDs in coordinate space defined via the LI construction ($\widetilde{\mathcal{H}}_{\rm LI}$) and the $\gamma_{0}$ operator ($\widetilde{\mathcal{H}}$). Upper: $P_{z}=1.937$ GeV, t=0.864-t=0.864 GeV2, lower: $P_{z}=1.937$ GeV, t=2.748-t=2.748 GeV2.
@@ III.3 Renormalized quasi-GPD Matrix Elements
To obtain physically meaningful results from lattice QCD calculations, the bare matrix elements extracted from correlation functions must be properly renormalized. In this work, we adopt the ratio scheme for renormalization as shown in Eq. (9). Fig. 4 shows the renormalized matrix elements $\widetilde{\mathcal{H}}^{R}$ as a function of the dimensionless variable $\lambda\equiv zP_{z}$ for four representative nonzero values of $\xi$. In each panel at fixed $\xi$, the magnitude of the matrix element decreases monotonically with increasing $|\lambda|$ for all momentum transfers, as expected from the suppression of long-distance contributions in the coordinate-space correlator. Additionally, at fixed $\lambda$ within each panel, we observe a systematic falloff in the matrix element magnitude as the momentum transfer t-t increases, consistent with the generic $t$-slope seen in hadronic form factors. These trends are consistent with short-distance dominance at small $z^{2}$ and provide stable inputs for the polynomiality-constrained fits employed in our moment extraction.
Figure 4: Renormalized matrix elements $\widetilde{\mathcal{H}}^{R}$ as functions of $\lambda$ at four nonzero skewness values. In each panel, data points are distinguished by color, representing results at different momentum transfers t-t.
@@ IV.1 Fitting strategy
For the fitting range, we fix the lower bound at $z_{\min}=0.08$ fm and vary the upper bound $z_{\max}$ to determine an optimal interval. Since the short-distance factorization is applicable only at short distances, we restrict $z_{\max}$ to the range $0.20\text{--}0.44~\mathrm{fm}$ and compare results obtained with LL, NLL, and NNLL resummation. Fig. 5 illustrates an example of the extracted moments at momentum $P_{z}=1.453~\mathrm{GeV}$ and momentum transfer $t=0.446~\mathrm{GeV}^{2}$, with the renormalization scale fixed at $\mu=2~\mathrm{GeV}$. We vary $\kappa$ among $1$, $\sqrt{2}$, and $2$, which are still close to one, to estimate the associated uncertainty. This range was also chosen after a study of the short-distance behavior of the same zero-momentum matrix elements in Ref. [88]. For the lowest moment $A_{1,0}$, RGR effects are negligible because it is associated with the conserved vector current, whose anomalous dimension vanishes. This behavior is indeed observed in the first panel, where the fixed–order and resummed results are essentially indistinguishable. In contrast, $A_{3,0}$ has a nonzero anomalous dimension, so logarithmic $\ln(z^{2}\mu^{2})$ terms appear in the matching, leading to visible sensitivity to the resummation as the fit range $z_{\max}$ is varied. At larger $z_{\max}$, the physical scale $\mu_{0}={(2\kappa e^{-\gamma_{E}})}/{|z|}$ decreases, driving the running coupling $\alpha_{s}(\mu_{0})$ into the nonperturbative regime and thereby spoiling the reliability of perturbation theory. From the $A_{3,0}$ panels, we observe that the LL-resummed result (matching coefficients at tree level) lies systematically below the NLL– and NNLL–resummed ones, illustrating the impact of perturbative matching. Moreover, varying $\kappa=\{1,\sqrt{2},2\}$ shows that the plateau region becomes progressively broader as $\kappa$ increases. This is mainly because a larger $\kappa$ allows for a wider range of $z$ to be included in the fit, while the weight is mainly carried by the short-$z$ matrix elements. To ensure stability and control systematic uncertainties, we therefore restrict the analysis to a short-distance region where fixed order (up to NNLO), NLL-, and NNLL-resummed results agree within the statistical uncertainties. Therefore, this is a definitive choice based on the criteria given, so it is not conservative; we adopt $z_{\max}=0.24~\mathrm{fm}$ in the following analysis. We further remark that when $\kappa\sim 1$ the nonperturbative effects are still sizable and do not provide meaningful constraints on $A_{5,0}$. In particular, no clear plateau is observed at $\kappa=1$, while the NNLL results show visibly better convergence than NLL. Accordingly, our choice of $z_{\max}$ is primarily informed by the plateau behavior observed in $A_{3,0}$.